xx Vectors

Vectors

Erik Schnetter <schnetter@cct.lsu.edu>

November 24 2010

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 testxall 10 Scope: private01  BOOLEAN

11 Description: test all available vector engines

11 Range 10              01  Default: no
11 yes  10 test all engines supported by the current CPU
11 no  10 test only default engine for current CPU

11

11 verbose10 Scope: private01  BOOLEAN

11 Description: Verbose output during vector tests

11        10              01  Default: no

11

7 Interfaces

General Implements: vectors

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 allxpassed 10 allxpassed 01 compact 11 0
11                    10                    01 description 11 all engines passed their tests
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT

Adds header: vectors.h

8 Schedule This section lists all the variables which are assigned storage by thorn CactusUtils/Vectors. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always: 11 x
11 allxpassed 11 x

11 x  11 x

11

Scheduled Functions

CCTKxSTARTUP x vectorsxstartup x print startup message x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPARAMCHECK x vectorsxtest x run correctness tests. x

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 allxpassed